# Learning What Reinforcement Learning Can’t: Interleaved Online Fine-Tuning for hardest Questions

### v1

@@ Abstract
Recent advances in large language model (LLM) reasoning have shown that sophisticated behaviors such as planning and self-reflection can emerge through reinforcement learning (RL). However, despite these successes, RL in its current form remains insufficient to induce capabilities that exceed the limitations of the base model, as it is primarily optimized based on existing knowledge of the model rather than facilitating the acquisition of new information. To address this limitation, we employ supervised fine-tuning (SFT) to learn what RL cannot, which enables the incorporation of new knowledge and reasoning patterns by leveraging high-quality demonstration data. We analyze the training dynamics of RL and SFT for LLM reasoning and find that RL excels at maintaining and improving performance on questions within the model’s original capabilities, while SFT is more effective at enabling progress on questions beyond the current scope of the model. Motivated by the complementary strengths of RL and SFT, we introduce a novel training approach, ReLIFT (Reinforcement Learning Interleaved with Online Fine-Tuning). In ReLIFT, the model is primarily trained using RL, but when it encounters challenging questions, high-quality solutions are collected for fine-tuning, and the training process alternates between RL and fine-tuning to enhance the model’s reasoning abilities. ReLIFT achieves an average improvement of over +5.2 points across five competition-level benchmarks and one out-of-distribution benchmark compared to other zero-RL models. Furthermore, we demonstrate that ReLIFT outperforms both RL and SFT while using only 13% of the detailed demonstration data, highlighting its scalability. These results provide compelling evidence that ReLIFT overcomes the fundamental limitations of RL and underscores the significant potential.
Project Page: https://github.com/TheRoadQaQ/ReLIFT
@@ 1 Introduction
Recent advancements in large language model (LLM) reasoning, as demonstrated by OpenAI-o1 Jaech et al. (2024), DeepSeek-R1 Guo et al. (2025), and Kimi-1.5 Team et al. (2025), highlight the potential to enhance reasoning capabilities by increasing test-time computational resources. These models typically generate long Chains-of-Thought (CoT Wei et al. (2022)) responses, exhibiting sophisticated behaviors such as reflection Kumar et al. (2024); Team et al. (2025) and planning Team et al. (2025). The primary factor driving this progress is large-scale Reinforcement Learning with Verifiable Rewards (RLVR), which does not depend on traditional scaffolding techniques like Monte Carlo Tree Search Xie et al. (2024); Li et al. (2025) or Process Reward Models Setlur et al. (2024). In RLVR, rewards are assigned based on whether the model’s output matches a ground-truth solution in mathematics or passes unit tests in code, thereby enabling scalability without the need for detailed demonstration data. In practice, RLVR is typically implemented using policy optimization algorithms such as Proximal Policy Optimization (PPO) Schulman et al. (2017) and Group Relative Policy Optimization (GRPO) Shao et al. (2024). Large-scale RLVR, characterized by its simplicity and directness, has enabled the emergence of long CoT reasoning and self-reflection capabilities.
@@
However, despite the empirical successes of RLVR, its current form is insufficient to incentivize capabilities that transcend the base model’s limitations Yue et al. (2025); Zhao et al. (2025). RL is inherently on-policy, learning from the LLM’s self-generated responses through iterative rollouts and feedbacks. Essentially, RL primarily enhances performance by biasing the model toward reasoning paths it already knows are more likely to yield rewards. In other words, most RL methods optimize based on the model’s existing knowledge rather than facilitating the acquisition of new information. In contrast, supervised fine-tuning (SFT) offers a complementary approach to enhancing the reasoning capabilities of LLMs. By leveraging high-quality demonstration data, SFT facilitates the incorporation of new knowledge and reasoning patterns Mecklenburg et al. (2024); Yue et al. (2025), and has been shown to achieve superior reasoning performance compared to RL, particularly in small-scale models Guo et al. (2025). Nevertheless, the effectiveness of SFT is highly contingent upon the availability of substantial amounts of high-quality demonstration data. Furthermore, models trained with SFT often exhibit limited generalization to out-of-distribution (OOD) scenarios, whereas RL-based approaches tend to demonstrate greater robustness Chu et al. (2025). These contrasting characteristics of RL and SFT highlight a promising research direction:
@@
To achieve this goal, we analyze the training dynamics by examining how the accuracy of the questions change under both RL and SFT. In Section 2.1, we evaluate the accuracy of the model on the questions at various checkpoints and classify these questions into four levels of difficulty. Our findings show that RL is more effective for relatively easy questions, whereas SFT is more beneficial for most difficult ones. Furthermore, for simple questions, SFT can degrade the model’s existing performance, as accuracy on these questions decreases and response length increases. In contrast, for challenging questions, the improvement provided by RL is less pronounced compared to SFT. Overall, RL efficiently trains the model to generate correct answers for questions it can already solve, while SFT is crucial for enabling the model to address questions that surpass its current capabilities.
@@
We conduct comprehensive experiments on five competition-level math reasoning benchmarks and one out-of-distribution (OOD) benchmark. Our proposed ReLIFT method achieves a new state-of-the-art accuracy of 51.1% based on Qwen2.5-Math-7B, outperforming previous zero-RL baselines by a significant margin, surpassing the strongest zero RL baseline by +5.2 points. Notably, ReLIFT demonstrates clear advantages over both pure SFT and pure RL, while requiring only 13% amount of detailed demonstration data. In addition, ReLIFT generates much more concise solutions, reducing average response length by nearly 10x compared to SFT, thus improving both performance and efficiency. We further validate the generality of ReLIFT by extending it to smaller and weaker foundation models, consistently observing superior results over standard RL and SFT approaches.
@@ 2 Reinforcement Learning Interleaved with Online Fine-Tuning
We start our study by examining how the model’s accuracy evolves on questions of varying difficulty during both SFT and RL training. Specifically, we independently perform SFT and RL on the Qwen2.5-Math-7B model, using a subset of the Open-R1-math-220K Face (2025). Both SFT and RL are conducted for 120 steps, with model checkpoints saved every 30 steps. Detailed experimental settings are provided in Appendix A.2. At each checkpoint, we evaluate model performance on a validation set comprising 1,000 questions. For a more detailed analysis, we categorize the validation questions into four distinct difficulty levels: Easy, Medium, Hard, and Hardest. To determine the difficulty level of each question , we prompt the Qwen2.5-Math-7B model with the same question eight times and calculate the average accuracy for that question, denoted as . Based on the value of , each question is assigned to a difficulty level according to the following criteria:
Easy: (i.e., 6, 7, or 8 correct out of 8)
Medium: (i.e., 3, 4, or 5 correct)
Hard: (i.e., 1 or 2 correct)
Hardest: (i.e., 0 correct)
Questions answered correctly at least 6 out of 8 times () are labeled Easy, as the model shows high consistency on these items. Questions never answered correctly () are labeled Hardest, exceeding the model’s capability. Intermediate thresholds ( for Medium and for Hard) are chosen to ensure each category has enough questions for reliable analysis. This classification allows a clear and systematic evaluation of model performance across difficulty levels, with similar numbers of questions in each group.
@@
We monitor the changes in average accuracy during RL and SFT training for Easy, Medium, Hard, and Hardest questions. As shown in Figure 1(a), RL outperforms SFT on Easy and Medium questions, while SFT yields better results than RL on Hardest questions. Notably, after SFT, the LLM fails to solve certain questions that were previously answered correctly by the initial model. In contrast, RL consistently improves accuracy across all difficulty levels and does not exhibit this issue. We also examine changes in the average response length during RL and SFT training, as shown in Figure 1(b). SFT produces longer responses across all question difficulties, aiming to mimic the target long response. In contrast, RL starts with shorter responses, and for medium and hard questions, response length increases over time, indicating more exploration. For the hardest questions, response length remains stable, suggesting the model stops exploring further. Ideally, responses to difficult questions should be longer, while those to simple questions should be shorter. However, SFT and RL each only partially achieve this objective.
@@
We further analyze the evolution of accuracy during training by tracking both the initial and final accuracy for each question, as illustrated in Figure 1(c). Our findings reveal substantial differences in how RL and SFT influence the model’s performance across questions of varying difficulty. For Easy and Medium questions, SFT results in decreased accuracy for some items, whereas RL generally preserves or enhances the model’s original ability to solve these questions. However, this pattern reverses for Hardest questions: SFT enables the model to learn and solve a greater number of these challenging items, as evidenced by the increased transition of Hardest questions into other categories. In contrast, RL does not yield comparable improvements for Hardest questions, highlighting a limitation in its ability to support learning beyond its current capabilities.
Based on these observations, we conclude that accuracy trajectories across questions of varying difficulty do not increase uniformly during RL and SFT training. Instead, a critical distinction emerges between these two paradigms: RL is more effective for Easy and Medium questions, whereas SFT proves more advantageous for the Hardest questions. These findings underscore the complementary strengths of RL and SFT: RL excels at maintaining and improving performance on questions within the model’s original capabilities, while SFT is more effective at facilitating progress on questions that extend beyond those capabilities. Consequently, by selectively incorporating the hardest questions during training and interleaving fine-tuning steps for these questions, it is possible to enhance the model’s overall performance and achieve more scalable and effective RL for LLMs.
@@ 2.2 Reinforcement Learning Interleaved with Online Fine-Tuning
Overview. Motivated by the complementary strengths of RL and SFT in addressing questions of varying difficulty, we propose a novel training approach termed Reinforcement Learning Interleaved with Online Fine-Tuning (ReLIFT). ReLIFT dynamically integrates RL and SFT based on the characteristics of the training data and the current capabilities of the model, as illustrated in Figure 2. During RL training, we identify particularly challenging examples based on rollout accuracy. For each of these hardest questions, we obtain high-quality CoT solutions either by collecting them in advance before training or by consulting a stronger model or human annotators, and filter out incorrect answers. These curated examples are added to an SFT buffer. When the buffer accumulates enough challenging questions to form a training batch, we perform an SFT step using these examples. The underlying intuition is that questions difficult for the model to learn through RL alone may benefit from targeted SFT. The timing of SFT is adaptive: when the model performs poorly on training questions, SFT is applied more frequently; as the model improves, RL is emphasized to further enhance its reasoning ability. Below, we provide a detailed description of our method.
@@
During the rollout phase of GRPO, for each question , we generate a set of solutions , which are evaluated by the reward function to obtain the accuracy . We identify particularly challenging questions as those for which . For these particularly hard questions, we solicit high-quality chain-of-thought (CoT) solutions from a stronger reasoning model (e.g., R1) or human experts, filter out incorrect CoTs to obtain reliable QA pairs , and add them to a buffer . This process is performed online during training and can be formalized as:
@@
Notably, the timing of interleaved fine-tuning is adaptive: at the early stage of training, when the model performs poorly on training questions, SFT is applied more frequently to help the model quickly learn better reasoning patterns. As training progresses and the model improves, RL is used more often to further incentivize and enhance the model’s existing reasoning abilities. Moreover, in most cases, it is unnecessary to prepare a large amount of CoTs in advance; instead, it is possible to only generate CoTs online for the hardest questions during training.
3 Experimens
3.1 Experimental Setting
Dataset Construction Our training set is a subset of OpenR1-Math-220k Face (2025), with prompts sourced from NuminaMath 1.5 Li et al. (2024) and detailed demonstrations generated by Deepseek-R1 Guo et al. (2025). We utilize the version of the dataset filtered by LUFFY Yan et al. (2025), resulting in 45,000 prompts and their corresponding CoT demonstrations.
RL and SFT Implementation. For RL, we omit the KL divergence loss term and set the entropy loss coefficient to 0.001. Consistent with Dr.GRPO Liu et al. (2025) and LUFFY Yan et al. (2025), we exclude both length normalization and standard error normalization from the GRPO loss 1 in all experiments. The rollout batch size is set to 128, the rollout times is 8, the update batch size is 64, and the learning rate is . A temperature of 1.0 is used during rollout generation. For SFT, we use a training batch size of 128 and a learning rate of .
ReLIFT Implementation. The RL and fine-tuning configurations for ReLIFT are the same as those described above. During interleaved fine-tuning, we utilize the demonstrations collected in advance. For RL and interleaved SFT, the learning rates are both set to .
Training. By default, we use Qwen2.5-Math-7B Yang et al. (2024), following previous works Yan et al. (2025); Cui et al. (2025); Liu et al. (2025); Zeng et al. ; Hu et al. (2025). To encourage greater exploration, and in line with LUFFY, we increase the RoPE theta from 10,000 to 40,000 and expand the window size to 16,384. We train for 500 steps for both RL and ReLIFT, and train 3 epochs for SFT. Further details are provided in Appendix A.1.
Baseline. (1) Simple-RL Zeng et al. : Trains Qwen2.5-Math-7B using a rule-based reward function. (2) Oat-Zero Liu et al. (2025): Also trains Qwen2.5-Math-7B with rule-based rewards, and further proposes removing the standard deviation in GRPO advantage calculation as well as omitting token-level normalization in the policy loss. (3) PRIME-Zero Cui et al. (2025): Utilizes policy rollouts and outcome labels, leveraging implicit process rewards. (4) OpenReasonerZero Hu et al. (2025): Trains Qwen-2.5-7B with rule-based rewards. (5) LUFFY Yan et al. (2025): Incorporates off-policy reasoning traces into RL by combining them with on-policy rollouts.

### v2

@@ Abstract
Recent advances in large language model (LLM) reasoning have shown that reasoning ability can emerge through reinforcement learning (RL). However, despite these successes, RL in its current form remains insufficient to induce capabilities that exceed the limitations of the base model, as it is primarily optimized based on existing knowledge of the model. To address this limitation, we employ supervised fine-tuning (SFT) to learn what RL cannot, which enables the incorporation of new knowledge and reasoning patterns by leveraging high-quality demonstration data. We analyze the training dynamics of RL and SFT for LLM reasoning and find that RL excels at improving performance on questions within the model’s original capabilities, while SFT is more effective at enabling progress on questions beyond the current scope of the model. Motivated by the complementary strengths of RL and SFT, we introduce ReLIFT (Reinforcement Learning Interleaved with Online Fine-Tuning), a novel training strategy. ReLIFT employs RL for general training, but interleaves it with targeted SFT on challenging questions for which high-quality solutions are collected online. By alternating between RL and SFT, ReLIFT addresses model weaknesses as they emerge. Empirically, ReLIFT outperforms previous RLVR methods by an average of +6.7 points across a suite of six benchmarks (five math reasoning and one out-of-distribution). More importantly, ReLIFT surpasses baselines such as individual RL, individual SFT, and various hybrid approaches while reducing the required training time. These results provide compelling evidence that ReLIFT is a powerful and resource-efficient paradigm for developing capable reasoning models. The code is available at here.
@@ 1 Introduction
Recent advancements in large language model reasoning, as demonstrated by OpenAI o-series Jaech et al. (2024), DeepSeek-R1 Guo et al. (2025), Gemini 2.5 Comanici et al. (2025), and Kimi k-series Team et al. (2025b), highlight the potential to enhance reasoning capabilities by increasing test-time computational resources. These models typically generate long Chains-of-Thought (CoT Wei et al. (2022)) responses, exhibiting sophisticated behaviors such as reflection Kumar et al. (2024) and planning Ke et al. (2025). The primary factor driving this progress is large-scale Reinforcement Learning with Verifiable Rewards (RLVR), which does not depend on traditional techniques like Monte Carlo Tree Search Xie et al. (2024); Li et al. (2025) or Process Reward Models Setlur et al. (2024); Zhang et al. (2025). In RLVR, rewards are assigned based on whether the model’s output matches a ground-truth solution in mathematics or passes unit tests in code, thereby enabling scalability without the need for demonstration data. Implemented with policy optimization algorithms like Proximal Policy Optimization (PPO) Schulman et al. (2017) and Group Relative Policy Optimization (GRPO) Shao et al. (2024), the simplicity and directness of large-scale RLVR have been pivotal in incentivizing the reasoning capabilities seen in today’s most advanced LLMs.
Despite its empirical successes, RLVR is insufficient to foster capabilities that transcend a base model’s inherent limitations Yue et al. (2025); Cheng et al. (2025); Zhao et al. (2025). Research suggests that RLVR primarily reinforces existing behaviors rather than instilling novel reasoning abilities. For instance, Yue et al. (2025) argue that RLVR fails to equip LLMs with new reasoning skills, noting that RLVR-trained models underperform base models on pass@k at higher k values, indicating limited expansion of reasoning capabilities. Similarly, Zhao et al. (2025) show that RL post-training amplifies certain pre-training behaviors while suppressing others. Furthermore, Cheng et al. (2025) find that RL stifles exploration, causing models to converge on narrow behavioral patterns and plateau on complex reasoning tasks. This limitation is rooted in the on-policy nature of RL, which learns from the model’s own generated responses via iterative rollouts and feedback. Consequently, RL primarily enhances performance by biasing the model toward reasoning paths it already knows are likely to yield rewards, optimizing existing knowledge rather than facilitating the acquisition of new information.
@@
In contrast, supervised fine-tuning (SFT) offers a complementary approach to enhancing the reasoning capabilities of LLMs. By leveraging high-quality demonstration data, SFT facilitates the incorporation of new knowledge and reasoning patterns Mecklenburg et al. (2024); Yue et al. (2025); Guo et al. (2025), and has been shown to achieve superior reasoning performance compared to RL, particularly in small-scale LLMs Guo et al. (2025). Nevertheless, the effectiveness of SFT relies highly on the availability of substantial amounts of high-quality demonstration data. Furthermore, SFT-trained models often exhibit limited generalization to out-of-distribution (OOD) scenarios, while RL-based approaches tend to demonstrate greater robustness Chu et al. (2025); Chen et al. (2025). These contrasting characteristics of RL and SFT highlight a promising research direction:
@@
To achieve this goal, we analyze the training dynamics of RL and SFT by examining how the accuracy of questions changes throughout the training process. We evaluate the model’s accuracy on questions at various checkpoints, and classify them into four levels of difficulty. Our findings indicate that RL is more effective for questions of lower difficulty, whereas SFT proves more beneficial for the most challenging ones. Furthermore, for simpler questions, SFT can degrade the model’s existing performance and the response length increases. In contrast, for challenging questions, the improvement provided by RL is less pronounced compared to SFT. Overall, RL efficiently trains the model to generate correct answers for questions it can already solve, while SFT is crucial for enabling the model to address questions that surpass its current capabilities.
@@
We conduct comprehensive experiments on five challenging math reasoning benchmarks and one OOD benchmark. Our proposed ReLIFT method achieves a new state-of-the-art accuracy of 52.6% with Qwen2.5-Math-7B, which significantly outperforms previous RLVR baselines. Notably, ReLIFT demonstrates clear superiority over methods including pure SFT, pure RL, and combined RL and SFT approaches, such as RL with SFT loss, SFT followed by RL, and LUFFY Yan et al. (2025). ReLIFT requires minimal detailed demonstration data and GPU training hours. Additionally, ReLIFT generates significantly more concise solutions, which improves both performance and efficiency. We further validate the generality of ReLIFT by extending it to smaller and weaker base models, consistently observing superior results.
The main contributions of this work are as follows:
We present a systematic analysis of the training dynamics of RL and SFT in reasoning, empirically demonstrating their complementary roles: RL refines existing skills on solvable problems, while SFT is essential for acquiring new knowledge for difficult ones.
We propose ReLIFT, a novel framework that interleaves RL with online fine-tuning. ReLIFT dynamically identifies challenging examples during RL for targeted fine-tuning to address emerging model weaknesses.
Extensive experiments show ReLIFT achieves state-of-the-art performance on mathematical reasoning and OOD benchmarks, outperforming pure SFT, pure RL, and existing hybrid methods with significantly less demonstration data and computational overhead.
@@ 2 Reinforcement Learning Interleaved with Online Fine-Tuning
We begin our study by examining how the model’s accuracy evolves on questions of varying difficulty during both SFT and RL training. Specifically, we independently train the Qwen2.5-Math-7B model using both SFT and RL on a subset of the Open-R1-math-220K dataset Face (2025). Both SFT and RL are conducted for 120 steps, with model checkpoints saved every 30 steps. Detailed experimental settings are provided in Appendix B. At each checkpoint, we evaluate model performance on a validation set comprising 1,000 questions. To facilitate a more detailed analysis, we categorize the validation questions into four distinct difficulty levels: Easy, Medium, Hard, and Hardest. To determine the difficulty level of each question , we prompt the Qwen2.5-Math-7B model eight times and calculate the average accuracy for that question, denoted as . Following the approach in works like Wen et al. (2025); Bae et al. (2025); Yu et al. (2025), and to ensure a sufficient number of questions in each category, we select these specific thresholds to enable a robust analysis of the training dynamics of both RL and SFT. The criteria for assigning difficulty levels are as follows:
Easy: 6 to 8 correct answers out of 8 ().
Medium: 3 to 5 correct answers out of 8 ().
Hard: 1 or 2 correct answers out of 8 ().
Hardest: 0 correct answers out of 8 (), representing items that are beyond the model’s current capabilities.
@@
We analyze the average accuracy of LLM trained with RL and SFT on four difficulty levels of questions: Easy, Medium, Hard, and Hardest. As shown in Figure 1(a), RL outperforms SFT on Easy and Medium questions. Conversely, SFT achieves better results than RL on the Hardest questions. Notably, after SFT, the LLM fails to solve certain questions that the initial model previously answered correctly. In contrast, RL consistently improves accuracy across all difficulty levels and does not exhibit this issue. We also examine changes in the average response length during RL and SFT training, as shown in Figure 1(b). SFT produces longer responses across all question difficulties, aiming to mimic the target long response. In contrast, RL begins with shorter responses, and for Medium and Hard questions, the response length increases over time, indicating more exploration. For the Hardest questions, the response length remains stable, which suggests the model stops exploring further. Ideally, responses to difficult questions should be longer, and responses to simple questions should be shorter. However, neither SFT nor RL fully achieves this objective.
@@
We further analyze the evolution of accuracy during training by tracking both the initial and final accuracy for each question, as illustrated in Figure 1(c). Our findings reveal substantial differences in how RL and SFT influence the model’s performance across questions of varying difficulty. For Easy and Medium questions, SFT results in decreased accuracy for some items, whereas RL generally preserves or enhances the LLM’s original ability to solve these questions. However, this pattern reverses for Hardest questions: SFT enables the model to learn and solve a greater number of these challenging items, as evidenced by the increased transition of Hardest questions into other categories. In contrast, RL does not yield comparable improvements for Hardest questions, highlighting a limitation in its ability to support learning beyond its current capabilities.
These findings demonstrate that RL and SFT have complementary strengths. RL excels at refining and preserving the model’s existing knowledge, making it suitable for questions within its competence. However, SFT is superior for acquiring new knowledge to tackle problems that lie beyond the model’s initial capabilities. This suggests that accuracy does not improve uniformly and that a hybrid training strategy could be optimal. By selectively interleaving SFT steps for the Hardest questions, it may be possible to enhance overall model performance and develop more scalable and effective RL for LLMs.
@@ 2.2 Reinforcement Learning Interleaved with Online Fine-Tuning
Overview. Motivated by the complementary strengths of RL and SFT in addressing questions of varying difficulty, we propose a novel training approach termed Reinforcement Learning Interleaved with Online Fine-Tuning (ReLIFT). ReLIFT dynamically integrates RL and fine-tuning (FT) based on the characteristics of the training data and the current capabilities of the model, as illustrated in Figure 2. During RL training, we identify hardest examples based on rollout accuracy. For each of these hardest questions, we obtain high-quality CoT solutions either by collecting them in advance before training or by consulting a stronger model or human annotators. These curated examples are added to a fine-tuning buffer. When the buffer accumulates enough questions to form a training batch, we perform a fine-tuning step using these examples. The underlying intuition is that questions difficult for the model to learn through RL alone may benefit from targeted fine-tuning. Below, we provide a detailed description of our method.
@@
During the rollout phase of GRPO, for each question , we generate a set of solutions , which are evaluated by the reward function to obtain the accuracy . We identify particularly challenging questions as those for which . For these particularly hard questions, we solicit high-quality CoT solutions from a stronger reasoning model (e.g., DeepSeek-R1) or human experts, filter out incorrect CoTs to obtain reliable QA pairs , and add them to a buffer . The data collected from is then added to a primary fine-tuning buffer, , which serves as the data for subsequent fine-tuning step. This process is performed online during training. Collecting and training can be done together, and can be formalized as:
@@
To prevent fine-tuning updates from overly constraining the model’s exploratory behavior, we incorporate an entropy regularization term into the loss function. The full objective is as follows:
where . The term represents the entropy of the token generation distribution , and is a hyperparameter controlling the weight of the entropy term.
@@
The frequency of this interleaved fine-tuning is adaptive. Early in training, when the model’s performance is low, SFT is applied more frequently to accelerate the learning of effective reasoning patterns. As training progresses and the model improves, RL is prioritized to further incentivize the model’s existing abilities. Furthermore, this online approach obviates the need for a large, pre-collected dataset of CoT demonstrations . Instead, CoTs can be generated dynamically only for the most challenging questions encountered during training.
3 Experimental Setup
Dataset Construction. Our training set is a subset of OpenR1-Math-220k Face (2025), with prompts sourced from NuminaMath 1.5 Li et al. (2024) and detailed demonstration data generated by Deepseek-R1 Guo et al. (2025). We use the default 94k prompt subset and filter out generations that are longer than 8192 tokens or are verified incorrect by Math-verify Face (2024). This process yields 46k prompts and high-quality demonstrations, which we term OpenR1-Math-46k-8192.